=1 manjubithe, mongGolbicig tailateh

manju bithe, mongol bicig-täi lateh^{1} Kirill "Us"agt"a"i LaTEX

Mongolian and Manju for L^{A}TEX 2ε

Implementation Level v.

Oliver Corff^{2}  Dorjpalam Dorj^{3}

1Instead of the transliterated text, the same text should appear in vertical Mongolian writing. Since did not find PostScript support on your system, it decided to print this text horizontally and in transliteration. It is suggested that you update your L^{A}TEX system by adding PostScript support (see section 2.4).

2The author received a DFG (Deutsche Forschungsgemeinschaft) postdoctoral grant (Habilitationsstipendium) while developing the first version of and expresses his sincere gratitude.

3Institute for Language and Literature, Mongolian Academy of Sciences, Ulaanbaatar.

Abstract

is now released in Implementation Level implying that there is not only Cyrillic Support for Modern Mongolian (Xalx dialect), Buryat and Russian; this version also includes comprehensive support for Mongolian Script (also known as Uighur Script) and Manju. All writings can be mixed freely within the same document, and within the same page.

The present release () is still very early; expect bugs and irregular behaviour. Especially the Mongolian full page mode still shows occasional quirks, depending on the document class loaded.

Contents

1

List of Tables

2

List of Figures

3

Part I

Chapter 1

: System Overview How to read and use the documentation

According to your specific needs, you can read various parts of this documentation separately.

First-time users interested in generating Cyrillic-only documents can confine themselves to the sections beginning on part II, page 21.

First-time users primarily interested in creating Mongol Bicig and Manju documents or text insertions in these languages can directly jump to section III, page 34.

Advanced users interested in improved Mongolian and Manju display functions should directly go to section ??, page ??.

Advanced users already familiar with the majority of functions can refer to the Alphabetic Command Reference beginning on page ??.

Nota Bene: Please note that includes the complete functionality of ManjuTEX which is hereby declared obsolete. ManjuTEX will not longer be supported.

System Features

2.1 Scope of

is a package which offers support for writing documents in Mongolian, Manju, Buryat and Russian. Mongolian can be represented in traditional Uighur script (also known as Classical or Traditional Script) and Cyrillic. Manju resembles the Traditional Mongolian script (from which it is derived) but uses a rich choice of diacritics in order to eliminate numerous ambiguities of the Mongolian script ancestor. Modern Buryat, like Mongolian in its present form, is written with a Cyrillic alphabet, but both Mongolian (35 letters) and Buryat (36 letters) use more letters than Russian (33 letters).

2.1.1 Mongolian

The word Mongolian is actually an umbrella term for several languages rather than the precise name of a single language. Things become more complicated when names of ethnic groups, languages and writing systems are mixed.

Xalx or Khalkha is the name of the Mongolian nationality residing in Mongolia proper. Their dialect forms the basis of Mongolian written with Cyrillic letters. Throughout this text, Modern Mongolian is used as a synonym.

Buryat is the name of the Mongolian nationality residing in Buryatia, north of Mongolia, east of Lake Baikal, being a part of the Russian Federation. The Buryat call themselves Buryaad while Xalx Mongolians call them Buriad. The English name follows the Russian orthography. Linguistically, Xalx and Buryat Mongol are fairly close languages; Buryat has a slightly different sound system in which the phoneme /s/ partially shifted to /h/; the modern Buryat Cyrillic alphabet (virtually identical with the Cyrillic alphabet used for writing Modern Mongolian) has one additional letter (H/h, H/h) for marking the difference to /s/.

Bicig (literally script in Mongolian) denotes text written in the traditional Mongolian script which is also referred to as Uighur. Throughout this document, the term Bicig will be used on an equal footing with Classical and Traditional Mongolian. The latter term is used in the name of the Unicode/ISO10646 character plane U1800 which contains Mongolian, Manju, Sibe and sets of special characters called Ali Gali or Galig. In order to identify Mongolian script related commands distinct for Mongolian and Manju, the Mongolian commands have the name root bicig whereas the Manju commands have the name root

     bithe.

Xalx Mongolian, or Modern Colloquial Mongolian, is about as different from the form written in Classical script as modern English in phonetical spelling (assume it be written in Shavian letters) from the highly historical orthography of Standard English. Beyond these differences, Mongolian written in Classical Script usually preserves a substantial amount of historical grammatical features which make it look a bit like Elizabethan English.

2.1.2 Manju

Manju is a Tungusic language closely related to Mongolian. Though Manju is virtually not spoken anymore, it has been the official language during 300 years of Manju government in Qing Dynasty China. Vast amounts of official documents survive, as well as some of the finest multilingual dictionaries ever compiled, e. g. the Pentaglot, or Mirror in Five Languages, a dictionary with 18671 entries in five languages (Manju, Tibetan, Mongolian, Uighur and Chinese). Manju writing is derived from Uighur Mongolian by adding diacritics in the form of dots and circles.

2.2 Implementation Levels

During several years of developing , the desired capabilities of a software package serving the described scope were classified and implemented along four Implementation Levels which have the following, well-defined properties:

Implementation Features

Level

I Modern Mongolian in Cyrillic Script and Buryat
II I and Mongolian script in LR mode horizontal
IIa II and Mongolian script portions in LR mode vertical
III II and Mongolian script text in horizontal RL mode
IV III and Mongolian script text printed vertically

Implementation Level I is good for producing documents in Modern Mongolian; Implementation Level II adds support for words and lines of Classical Mongolian embedded in other (Cyrillic and Latin) scripts which is essentially useful for dictionaries etc.; Implementation Level IIa allows single words to be placed in vertical capsules; Implementation Level III allows the composition of purely Classical Mongolian documents while Implementation Level IV finally allows the combination of both scripts in freely assignable quantities and locations within the document.

Mongolian linguistic culture provides a perplexing richness of writing systems of varying regional, historical and socio-political importance. Developing a Mongolian system which covers all documented writing systems is tantamount to writing a Mongolian Babel system and cannot be done in a truly elegant manner with respect to the current L^{A}TEX 2ε limitations.

2.3 Requirements and Limitations

In order to run a recent version of L^{A}TEX 2ε is necessary. relies on the NFSS font selection mechanism and the ligature capacities of Metafont. This package has not been tested under L^{A}TEX2.09 and will most certainly not function satisfactorily under that environment. Depending on the implementation level, further software support becomes necessary since not all features can be realized smoothly in L^{A}TEX 2ε alone.

Implementation Level Requirements
I A working L^{A}TEX 2ε system
II A working L^{A}TEX 2ε system
IIa like II, plus PostScript support
III like II, plus functional TeX–XeT system
IV like III, plus PostScript support

The eTEX (available for DOS and UNIX based computers alike) system provides full Right-to-Left writing support; eTEX and eL^{A}TEX 2ε are part of all modern TEX implementations for the majority of operating systems^{1}.

PostScript support is a standard feature of most UNIX installations and is also supplied with most of the available TEX for Windows distributions^{2}.

A word of warning is necessary here. is not a small, convenient system which can be used without any effort. Much like its very foundation L^{A}TEX it requires some willingness to study a few (and indeed simple) rules; occasionally one or the other old habit has to be overcome. The reward is text typeset in excellent quality so that scholarly achievement no longer disappears in badly typeset documents.

1It is also possible to build an eTEX system from scratch using the web2c (or teTeX)

sources, replacing tex.web with tex–xet.web and tex.ch with tex–xet.ch.

2The authors used Linux (Red Hat 4.2 through 7.1) systems for the developing work; on the same hardware, PostScript under Windowsxx is significantly slower than under Linux; this holds true for document and font compilation as well.

$ elatex montex.tex     # Compile document
$ dvips montex.dvi      # Create PostScript out of DVI
$ gv montex.ps          # Preview document

Figure 2.1: PostScript Compilation and Preview Cycle

2.4 PostScript Support

PostScript is used for creating vertical capsules of text within horizontal text for implementation levels IIa as well as for complete pages with implementation level IV. This requires the presence of the rotating package for L^{A}TEX which itself relies on the graphics package. The rotating and graphics packages come with teTeX but do not come with emtex. They can be found at CTAN.

Besides the above-mentioned packages it is necessary that the generated .dvi files can be processed further, e. g. by dvips which generates a PostScript file out of a .dvi file. If there is no PostScript printer at your site, PostScript emulation is necessary which is usually provided by GhostScript and GhostView. Implementations of these systems are available for a large number of operating systems and can also be found at CTAN. Linux, a free UNIX system, comes which GhostScript, and the winemtex distribution of L^{A}TEX includes GhostScript as well. A sample command sequence to produce and preview a document with these utilities can be found in illustration 2.1.

Without PostScript support, only implementation level II can be realized (instead of IIa and IV): Mongolian script can be printed horizontally but not vertically. It must be noted here that most DVI viewers are not capable of presenting vertical text correctly; the conversion step from DVI to PostScript is virtually always necessary.

2.5 PDF Support

With the arrival of PDFTEX it is possible to generate a PDF (Adobe’s Portable Document Format, that is) directly from the .tex sources without going through the .dvi stage. All systems offering pdfelatex can be used to compile a Implementation Level IV document. provided the necessary PostScript is installed. PDF is the recommended form of output on systems without PostScript views and printers (like, unfortunately, most of the Windowsxx world). PDF documents reproduce everything generates as is, and with Type1 fonts for Classical Mongolian, the on-screen display of Classical Mongolian and Manju material is fast and pleasant.

A sample command sequence to produce and preview a document with PDF output can be found in illustration 2.2.

It is recommended to users with menu-driven environments (WinEdt,

$ pdfelatex montex.tex     # Compile document
$ acroread montex.pdf      # View PDF Document with Acrobat Reader

Figure 2.2: PDF Compilation and Preview Cycle

TeXnicCenter, TeXshell etc.) to set the compilation commands to the effect that pdfelatex is invoked as the default compilation engine, and the Acrobat Reader is invoked as the default viewer. Please consult the software documentation of these products for the necessary steps and procedures.

Acknowledgements

The authors wish to thank the creators of TEX and L^{A}TEX as well as the designers of the existing fonts for their generosity of providing the world with such inspiring pieces of software. The packages from which pieces of code originated by inspiration or blunt copy are far too numerous; the Russian captions were taken from the file russian.sty (as were the English captions), most of the Cyrillic letters were produced with fonts by Nana Glonti and Alexander Samarin; additional letters were taken from J. Knappen’s font files. Special thanks go to David Carlisle who offered the solution for a serious problem with the ligature mechanism in TEX. During the development of Implementation Level IV, important suggestions came from David Kastrup, Robin Fairbairns, Dan Luecking, e. a. Intensive communications about Cyrillic fonts and integrating with the LH fonts took place with Vladimir Volovich, and other problems were discussed with Werner Lemberg.

Among the friends and colleagues in Mongolia and Germany who offered information, support and encouragement the authors wish to name B. Nerguy, Urgamal, M. Balk, Q. Änxzayaa and K. Maezono (without implication of any particular order or precedence). They contributed test data as well as their ideas for encoding, font shapes, user interfaces, and, last but not least, were patient alpha testers who helped the authors with numerous problem reports.

Many of the improvements between version 0.1 and the present version are not actually improvements; they are simply eliminations of partially awful bugs as well as ugly hacks (rather than code) and aim to make this package simply usable (if not useful).

3.1 Sources of Code and Inspiration

Some Cyrillic packages have been available for a few years. All Cyrillic packages available for TEX/L^{A}TEX stem from one of two lines of ancestry:

• Fonts developed at the University of Washington

• Fonts by Nana Glonti and Alexander Samarin

The two lines differ substantially in scope of characters and printing quality. The University of Washington series in OT2 encoding has a broad support for East European languages, but the praise for printing quality is given to the characters designed by Nana Glonti and Alexander Samarin. The Glonti/Samarin line of characters has undergone numerous minor modifications, not so much in glyph shapes but basically in determining encoding slots. Fortunately, the fonts are set up in a way that allows for convenient redefining of individual code positions.

Only in 1999, a comprehensive set of Cyrillic glyphs in various encodings called LH was finally implemented as standard Cyrillic support for L^{A}TEX 2ε, but at that time it was decided that for the time being will continue to offer its own Cyrillic font set, for which there are mainly three reasons:

1. The set has a seven-bit basis and allows for the all-Latin, all-ASCII communication of Mongolian texts, while Mongolian hyphenation is active—a matching LH encoding is not yet established;

2. offers a Mongolian currency sign which is not yet included in the LH fonts;

3. Mongolia, on of the prime markets for , continues to use partially outdated L^{A}TEX 2ε installations.

Neither original line of Cyrillic characters offers the additional characters necessary for writing non-Slavic languages like Mongolian; already in the beginning of the 1990s, Jörg Knappen filled the gap and designed additional letters which were intended to be used with Bashkir, a Turkic language. In fact, most of the letter forms employed there can also be used in other non-Slavic languages used throughout Central Asia since these letters are not specific to Bashkir. Some of J. Knappen’s letter forms (accidentally mostly those which are not necessary for writing Modern Mongolian) do need some refinement, and are then immediately suitable for a range of languages including Kasakh, Tuvinian etc. In the present stage of the system, only those letters used in Mongolian and Buryat are incorporated from J. Knappen’s files.

After discussing the typeface issues with Mongolian specialists, the Glonti/Samarin letter forms were chosen for their superior appearance in volume text. The fonts had to be renamed; failing to do so would have resulted in unpermissible ambiguity.

One feature of the traditional Cyrillic font packages for TEX (besides their lacking support for non-Slavic languages) is the intimate relationship between input encoding and output encoding. The first step in building Mongolian support was to separate these two spheres as numerous Mongolian encodings exist which should all be supported by the Mongolian package. A new encoding was then defined ( — Local Mongolian Cyrillic) which is a close approximation of a transliteration based on Latin1 encoding, notably with front vowels ä, ö, ü ("a, "o, "u) and ï ("i) in matching positions.

The encoding is completely detached from the existing Cyrillic codepages of which there are too many; in addition it should be possible to produce Mongolian documents in 7-bit environments so as to assure maximum document portability.

An additional ligature table for Metafont was then supplied which takes care of most of the two-letter combinations necessary for entering Cyrillic since the Cyrillic alphabet has more letters (36 in the present version) than the Latin alphabet which prohibited any 1:1-mapping scheme.

The used transliteration is very closely modelled after the MLS system yet provides enough transparency for accepting alternative spellings in some cases.

Input and Output Encodings

4.1 The Need for Encodings

Any Mongolian text system has to deal with the issues of how to store, transmit, process and represent the following entities:

• Normal Latin letters, numbers and punctuation marks: a, b, c, etc.;

• Cyrillic letters, including those not present in basic Cyrillic but needed by Mongolian: a, b, w, ö, ü;

• Special symbols like the Mongolian Currency sign: ;

• Classical Mongolian script;

• Special symbols used in Latin scripts for purposes of transliterating Mongolian scripts: ä, ö, ü, e. a.

All these sets of symbols, alphabets and characters have their own unique properties, especially when it comes to non-Latin writings like Mongolian or Tibetan.

Unfortunately, prior to the arrival of Unicode, all computer systems based on 8-bit encodings (with a maximum character set of 256 characters) can only represent subsets of the above-mentioned entities as basic characters. All computer systems with 8-bit character encodings must either switch between several character sets (or code pages) or use non-standard commands to invoke individual character entities.

It is important to understand that the issue of how to enter all these characters is more or less completely detached from the issue of how to represent these characters on screen or in a document. Misleadingly, the usage of Latin characters in modern computers seems to suggest that there is a simple, 1:1 relationship (or mapping) between input and output, but for a number of languages, including Mongolian, this is simply not the case. While, due to the origin and history of computers, simple-minded systems do not make any difference between the two realms, TEX separates these two domains clearly, allowing for the amazing flexibility TEX shows when treating languages and writing systems.

It must also be understood that even though Unicode allows for the unambiguous representation of the characters and symbols of the world’s major languages, it does not define any output conventions, and thus, input and output domains should still be treated as separate areas.

4.2 Input Encodings

is flexible enough to deal with several kinds of input encodings including code pages with Cyrillic letters and Unicode. Input encodings are declared as an option to the main package in the document preample. E. g., a user working on an IBM compatible DOS platform is likely to specify the option mls:

\usepackage[mls]{mls}

4.2.1 7-bit ASCII and Mongolian

Basically it is possible to use without anything else but the plain 7-bit ASCII Latin character set since internal and external mechanisms are available which can render transliterated texts (both Cyrillic and Traditional Mongolian) into their appropriate script presentations.

4.2.2 The MLS Codepage

The MLS codepage was the ancestor of all comprehensive, IBM-compatible Mongolian systems which intended to cover both Cyrillic and Classical Mongolian. Developed in the early 1990s, the MLS system tried to offer full Mongolian support for existing hardware and software as it was available then. While modern technological developments have confined the original approach to history, it is preserved here for preserving backward compatibility.

The MLS codepage is compatible with the IBM 437 codepage as far as the front vowels are concerned but features additional Cyrillic letters and Classical Mongolian.

4.2.3 8-bit Encodings

Most available 8-bit input encodings support either front vowels or Cyrillic letters or Classical Mongolian but usually not several of them at the same time.

If a local environment supports Cyrillic and Script codepages then texts can be composed using these codepages. Table 4.1 shows which codepages are supported. Those codepage names which are followed by a ‘(+)’ are supplied by whereas the other codepage declarations are recognized and passed through to the system assuming that the appropriate table exists. The column “Front Vowels” indicates whether the vowels ö and ü (and their Mongolian counterparts "o, "u) are available in that particular codepage. recognizes both numbers and numbers preceded by cp, like 1250 and cp1250 as names of codepages which are known by their number.

4.2.4 utf-8 Unicode

In summer of 2002, a new input encoding was made available for existing L^{A}TEX 2ε installations which allows the processing of utf8-encoded Unicode material. This package can be invoked with the option utf8:

\usepackage[utf8]{mls}

There are some caveats, however. The relevant code is still under development, and at present, only deals with the Mongolian, Manju and Sibe subsets of the Traditional Mongolian Character Plane beginning at U1800; a Todo character set remains to be implemented, and some of the more arcane special characters present in Unicode are as yet unavailable in . The resulting constraints do not affect the work with contemporary text material and are only felt when dealing with frequently bilingual, mostly Tibetan and Sanskrit, religious texts of earlier centuries. Consult chapter ?? on page ?? and table ?? for details.

4.3 Output Encodings

Several output encodings are defined for :

Local Mongolian Cyrillic: This encoding was defined in order to avoid collisions with existing Cyrillic encodings for TEX and L^{A}TEX. is a 7-bit encoding which implies that most of its characters are addressed in the range of ordinary ASCII characters; when this encoding is active, all text typed in ASCII Latin characters will automatically appear in Cyrillic. Unlike some other available 7-bit encodings (like WN Cyrillic) it provides characters used in Mongolian.

Local Manju: Manju in Ligature Mode. Any text typed in romanized Manju is automatically converted into Manju characters. acts thus like a typical 7-bit encoding.

Local Mongolian: Similar to Manju in Ligature Mode, Mongolian in Ligature Mode is typed in a special romanized form and is then automatically converted into Uighur Mongolian characters. , too, acts thus like a typical 7-bit encoding.

Enc. Option  Latin Transliteration Symbols  Cyrillic Front Vowels
Ä/ä Ö/ö Ü/ü Ë/ë Ï/ï Č/č Š/š  "O/"o "U/"u
mls (+)  +  +  +  ë  ï  -  -  +  +  +
ncc (+)  -  -  -  -  -  -  -  +  +  +
mos (+)  -  -  -  -  -  -  -  +  +  +
mnk (+)  -  -  -  -  -  -  -  +  +  +
dbk (+)  -  -  -  -  -  -  -  +  +  +
ctt (+)  -  -  -  -  -  -  -  +  +  +
ibmrus(+)  -  -  -  -  -  -  -  +  -  -
koi (+)  -  -  -  -  -  -  -  +  -  -
437  +  +  +  ë  ï  -  -  -  -  -
437de  +  +  +  ë  ï  -  -  -  -  -
850  +  +  +  + + -  -  -  -  -
852  +  +  +  + -  -  -  -  -  -
865  +  +  +  ë  ï  -  -  -  -  -
1250  Ä  +  ü  + -  -  -  -  -  -
1252  +  +  +  + + -  -  -  -  -
applemac  +  +  +  + + -  -  -  -  -
mac  +  +  +  + + -  -  -  -  -
ansinew  +  +  +  + + -  +  -  -  -
ascii  -  -  -  -  -  -  -  -  -  -
atari  +  +  +  + Ï  -  -  -  -  -
decmulti  +  +  +  + + -  -  -  -  -
isolatin  +  +  +  + Ï  -  -  -  -  -
latin1  +  +  +  + + -  -  -  -  -
latin2  +  +  +  + - + +  -  -  -
latin3  +  +  +  + + -  -  -  -  -
latin5  +  +  +  + + -  -  -  -  -
next  +  +  +  + + -  -  -  -  -
pc850  +  +  +  + Ï  -  -  -  -  -
roman8  +  +  +  + Ï  -  -  -  -  -

Table 4.1: Input Encodings Local Mongolian Script: The system’s original encoding for the Mongolian

script. Mongolian is represented by a Latin transliteration the letters of which are essentially treated as future Mongolian canonical code positions. Once Mongolian Unicode will be available, the Latin transliteration can be seamlessly replaced by Mongolian canonical characters. The arrangement of code positions in this encoding does not reflect Unicode but follows the MLS system’s keyboard support.

Local Mongolian-Tibetan: This encoding is reserved to ensure access to the characters in the future package: ^{1} and . It is designed to comprise Tibetan as well, and Sirlin’s Tibetan fonts can be directly used with this encoding.

Local Mongolian : This encoding is used for the Script (available on CTAN) but is not frozen yet. Individual code positions are still subject to change.

Local Mongolian Superset (U stands for ‘Umbrella”, “Unknown”, or whatever you like to pick): This encoding is used to access all glyphs of the bxg glyph container, but is not frozen yet. Individual code positions are still subject to change.

4.4 and Recent TEX Trends

As soon as the LH Cyrillic fonts support the Mongolian currency sign, will switch to this font set. At the moment the private encoding is favoured over LH; future implementations of will provide a smooth transition for the user: documents developed with older versions of will be upward compatible.

The babel package will, perhaps, also be supported in due course; at the moment, babel support is lacking mainly due to font encoding questions and a private RL setup. At present, is not built with babel compatibility in mind. It must be seen as a stand-alone extension similar to german.sty or the CJK package.

The future belongs to 16-bit character sets; the first TEX development supporting larger character sets is Ωmega of which experimental versions exist. One of the great features of Ωmega is the capability to process canonical input encodings in order to generate glyph variants for document presentation. These so-called translation processes are far more powerful than anything Metafont can offer via ligatures, and they are the only feasible way to avoid external preprocessors or internal retransliteration engines coded in TEX needed to process Mongolian script.^{2} Prof. Lagally’s ArabTEX is the only L^{A}TEX package known to the authors where an extensive retransliteration engine is realized as pure TEX code; it is an impressive piece of work defying any simple-minded imitation. So far, Ωmega translation processes exist for Tibetan and Arabic (paragons of complex relations between original script and any attempted romanization).

1It is possible to use the Soyombo package available since 1996 as long as is not available.

2The retransliteration engine provided with the encoding of has a rather ‘combined’ approach; basic letter forms are selected in the retransliteration section while typical ligatures are composed with the ligature tables of Metafont. The authors express their sincerest gratitude to David Carlisle who contributed the missing link between characters in the output list and TEX/Metafont’s ligature mechanism.

Installation

Before this latest version of is installed please make sure that old installations of and ManjuTEX are purged from disk as there are file name conflicts between earlier and recent versions of this software. In addition, ManjuTEX is not required any more as its functionality is now completely covered by .

5.1 Hyphenation Patterns

provides hyphenation rules for Modern Mongolian (Xalx). Hyphenation patterns for English are activated with English as selected language; hyphenation patterns for Russian exist at CTAN but they are unfortunately not suited for withour prior work. Hyphenation patterns for Buryat have not been developed yet.

Due to the very nature of TEX, hyphenation patterns for a given language cannot easily be loaded at run-time but must be compiled into a so-called format file which gets loaded by TEX whenever the command latex is executed. A format file is usually created when a new TEX or L^{A}TEX 2ε system is installed, but creating a new format can be done at any later time again. A special variant of TEX called initex is used for this purpose. The procedure sounds more intimidating than it actually is. Since there are many different types of TEX installations, the procedure is somewhat system-dependent. There is detailed on-line documentation available for performing this task, either in form of a text file for emtex, or in form of a FAQ file which can be displayed using the command texconfig faq on teTeX systems.

Part II

User Commands I General Settings Cyrillic Input Introduction

With regard to the substantial differences between Latin-like scripts (including Cyrillic) and Mongolian scripts, the user documentation of is divided into two parts. This part deals with general settings, like language choices and input encoding definitions, whereas the commands specific to Mongolian and Manju are dealt with in part III, “Mongol Bicig and Manju Bithe”. An alphabetic command reference covering all commands is presented in part IV.

6.1 General Settings

In order to access the commands of the package must be loaded in the document preamble by saying

\usepackage[<language options>,<encoding options>]{mls}

The options include choices for the basic document language and input encodings.

6.1.1 Document Language

The document language can be set with one of bicig, bithe, buryat,

english, russian or xalx like in
\usepackage[xalx]{mls}

which issues all captions and the date in Modern Mongolian.

The options bicig and bithe are discussed extensively in part III, “Mongol Bicig and Manju Bithe”.

The options buryat (see table ??), russian (see table ??) and xalx (see table ??) produce captions in Buryat, Russian and Modern Mongolian.

Buryat

Russian

The option english, at least as a \usepackage option, is essentially a do-nothing: it sets captions to English (which is the default of this package anyway).

The date form follows TEX conventions and is thus a mixture of numbers and words. Thus for today (July 4, 2002) we get^{1} what is shown in table 6.1. The Uighur Mongolian and Manju dates are presented in section 8.2, page 36.

Buryat Xalx Russian

Figure 6.1: Dates in Buryat, Xalx and Russian

The language specifiers buryat, english, russian and xalx can also be used anywhere in the document as arguments to the \selectlanguage command. Instead of stating an argument to \usepackage[...]{mls} it is possible to say in your document

     \selectlanguange{xalx}

which would set captions to Xalx Mongolian.

1The actual date at compilation time is used for the examples.

Xalx 24

Cyrillic Text – Kirill "us"ag

7.1 Cyrillic Text in Transliteration () Mode

provides two basic modes of operation: in

• Transliteration Mode (intimately linked to the encoding) all incoming text is regarded as transliterated Cyrillic. This allows users to compose Cyrillic documents on pure ASCII machines. In contrast, the

• Immediate Mode does nothing and waits for explicit Cyrillic characters in the input in order to generate Cyrillic output.

Two commands are used to switch between these modes:

     \SetDocumentEncodingLMC
     \SetDocumentEncodingNeutral

The first command switches to Transliteration Mode, the second command deactivates the transliteration and thus, by definition, activates Immediate Mode.

In the encoding, most Cyrillic characters are mapped directly to a single Latin character but for some characters there is a text command which became necessary since there are more Cyrillic than Latin characters. For convenience, a few ligatures were defined, too. Details are given in table 7.1.

Front vowels can be entered directly using the encoding slot of a valid and active input encoding, or they can be expressed via an abbreviated "v notation where v stands for any desired vowel. In the encoding used by , " is not an active character; selecting the proper letter is done by ligature statements in the Metafont sources.

Some letters can be entered with or without a preceding \, like and . Both \yu and yu will produce a . While yu is interpreted as a ligature, \yu allows for the character to be combined with accents. Accents are not commonly used in Mongolian since there are precise rules for word stress.

Cyrillic Letter  Input  Generic Command

 1 A  a  A      a    \CYRA      \cyra
 2 B  b  B      b    \CYRB      \cyrb
 3 W  w  W      w    \CYRV      \cyrw
 4 G  g  G      g    \CYRG      \cyrg
 5 D  d  D      d    \CYRD      \cyrd
 6 E  e  E      e    \CYRE      \cyre
 7  Ë/"E   ë/"e   \CYRYO     \cyryo
                  {\}YO {\}yo
 8 J  j  J      j    \CYRZH     \cyrzh
 9 Z  z  Z      z    \CYRZ      \cyrz
10 I  i  I      i    \CYRI      \cyri
11  Ï/"I   ï/"i   \CYRISHRT  \cyrishrt
                  {\}YI {\}yi
12 K  k  K      k    \CYRK      \cyrk
13 L  l  L      l    \CYRL      \cyrl
14 M  m  M      m    \CYRM      \cyrm
15 N  n  N      n    \CYRN      \cyrn
16 O  o  O      o    \CYRO      \cyro
17  Ö/"O   ö/"o   \CYROTLD   \cyrotld
18 P  p  P      p    \CYRP      \cyrp
19 R  r  R      r    \CYRR      \cyrr
20 S  s  S      s    \CYRS      \cyrs
21 T  t  T      t    \CYRT      \cyrt
22 U  u  U      u    \CYRU      \cyru
23 "U  "u  Ü/"U   ü/"u   \CYRY      \cyry
24 F  f  F      f    \CYRF      \cyrf
25 X  x  X      x    \CYRH      \cyrh
26 H  h  H      h    \CYRHSHA   cyrhsha
27 C  c  C      c    \CYRC      \cyrc
28 Q  q  Q      q    \CYRCH     \cyrch
                   \Ch    \ch
29  \Sh    \sh   \CYRSH     \cyrsh

                          sh

30  \Sc    \sc   \CYRSHCH   \cyrshch
                   \Qh    \qh
31  \Y     \y   \CYRHRDSN  \cyrhrdsn
32 Y  y  Y      y    \CYRERY    \cyrery
33  \I     \i   \CYRSFTSN  \cyrsftsn
34  Ä/"A   ä/"a   \CYREREV   \cyrerev
35 YU  yu  {\}YU {\}yu  \CYRYU     \cyryu
36 YA  ya  {\}YA {\}ya  \CYRYA     \cyrya

Table 7.1: Cyrillic Alphabet Input Methods This feature is taken from the OT2 encoding and is included mainly for the sake of completeness, convenience and compatibility^{1}.

Here now a sample of Mongolian text:

«Xalxyn gurwan "ond"or» x"am"a"an aldarshsan, Z"u"un xyazgaaryg toxinuulax sa"id N. Dugarjaw ardyn xuwısgalyn b"u"ur "ax"an "ue"as xamgi"in "agz"agt"a"i amı d"u"is"an alband tomilogdox c"ar"ag da"iny olon quxal daalgawryg xiq"a"ang"u"il"an biel"u"ulj yawsan t"u"uxt"a"i x"un.

{\mnr<<Xalxyn gurwan "ond"or>>
x"am"a"an aldarshsan, Z"u"un xyazgaaryg
toxinuulax sa"id N.~Dugarjaw ardyn
xuw\i sgalyn b"u"ur "ax"an "ue"as
xamgi"in "agz"agt"a"i am\i\ d"u"is"an
alband tomilogdox c"ar"ag da"iny olon
quxal daalgawryg xiq"a"ang"u"il"an
biel"u"ulj yawsan t"u"uxt"a"i x"un.}%

Figure 7.1: Romanized Cyrillic Input Example

In order to make the document you are reading at the moment truly portable, the somewhat more clumsy "v notation was used in this example; if your environments supports an 8-bit codepage (what it usually does), all front vowels can be entered as ä, ö and ü etc. using the slots of those vowels in the particular active codepage.^{2}

7.2 Entering Cyrillic Text in Immediate Mode

For freely combining Latin and Cyrillic characters without using any explicite commands it is necessary that the codepage in use supports some Cyrillic encoding. It should be noted, however that these documents are not easily portable between different platforms anymore since they need recoding; some of the Cyrillic codepages are defective in one or the other way thus individual characters can get lost.

The user simply specifies the desired input encoding as a \usepackage[<encoding>]{mls} option, and takes care of the rest. It is a feature and not a bug that input encoding and document language are chosen independently. It is well possible that a user working on a computer with default Mongolian codepage wants to create a document in Russian, English or any other language yet wants to include Mongolian fragments in her text without explicitely issuing any command.

1The magic triple-C!

2Looking at the source code of this document the astute reader will discover that all front vowels are indeed produced using the "a (etc.) notation; thus the document source can be viewn and manipulated on any 7-bit ASCII platform; it can also safely be transmitted via e-mail.

In case a need arises for switching from Transliteration Mode to Immediate Mode the command can be issued anywhere in the preamble or the document itself; like \SetDocumentEncodingLMC it affects the Cyrillic transliteration only and leaves the document language in its chosen state.

7.3 Entering Cyrillic Characters by Name

Outside the Cyrillic environments, individual Cyrillic characters can be entered by using the commands beginning with \cyrx from the two right columns of table 7.1 where x stands for the letter name. This command works in any encoding.

7.4 Entering Special Cyrillic Characters

A few special characters are available, notably the guillemots frequently used for quoting text, the currency symbol, the ordinal number symbol and the currency sign. See table 7.2.

Symbol Command Alternative

                         \lgu       <<
                         \rgu       >>

                         \No
                         \Togrog
                         \togrog

Table 7.2: Special Cyrillic Characters

The command producing the guillemots (\lgu, \rgu) only works in a Cyrillic environment — it is not a generic command.

There are actually two versions of the togrog command. While Togrog produces a sans serif (considered standard) with any font selected it is also possible to print serif (), italic () and typewriter () versions of this symbol.^{3} For achieving this result the commands MyTogrog and mytogrog are available. Unlike the standard command they simply pick the current font style of the surrounding letters for the currency symbol.

3The currency symbol is not limited to these three typefaces; all typefaces can be selected.

7.5 Running Text with Embedded Words in Different Encodings

Independently of the document language it is possible to produce portions of Cyrillic text within Latin text and vice versa. The two commands mnr and rnm switch from ordinary Latin text to transliterated Cyrillic text and back to Latin text. The command stands for mongolian new romanization and its reversal (which can, by accident, also be read as return to normal). They can be used as stream commands or for initializing groups:

mongol x"al ba english text with a mongol word inserted

\mnr mongol x"al ba \rnm english
text with a {\mnr mongol} word
inserted

For enhanced convenience, portions of text can also be encapsulated into the commands xalx{...} for Cyrillic text and lat{...} for neutral (i. e. Latin) texts.

   The commands \mnr, \rnm, \xalx{...} and \lat{...} do not switch

the default encoding; this shows up when a construct like \lat{\verb|article|} is placed in Transliteration Mode; the result will be article rather than article; in order to generate the desired form, the mode switching commands must be used.

7.6 Font Selection Commands

The Cyrillic fonts are set up in a manner which allows for seamless switching between Roman and Cyrillic typefaces. The font switching commands used for modifying typefaces (by \text..) are completely transparent to the encoding; no precaution whatsoever has to be taken. Most of the typefaces supplied with the traditional OT1 encoding are also available for ; Dunhill and Funny Roman are included.^{4} offers the following font families as shown in table 7.3:

The word “Roman” was avoided since in these families also cover matching typefaces in Cyrillic script. The first three families have support for combinations of different weights and shapes (e. g. bold and italic) whereas the other series usually only offer an italic variant. The Sans Serif Quotation Style 8pt typeface is not by default installed in standard L^{A}TEX distributions hence it cannot be guaranteed that switching to and from Cyrillic letters maintains the typeface. The fonts (upright and slanted) can be accessed via the \fontfamily{cmssq} command but are not shown in table 7.4. See table 7.4 for a therefore incomplete list of available typeface examples.

4A complete overview of the NFSS classification of the Computer Modern fonts can be found in The L^{A}TEX Companion, by Michel Goossens, Frank Mittelbach and Alexander

  \fontfamily{...}

Parameter

Family Description

cmr  Computer Modern Serif
cmss  Computer Modern Sans Serif
cmtt  computer Modern Typewriter
cmvtt  Computer Modern Variable Width Typewriter
cmfr  Computer Modern Funny
cmfib  Computer Modern Fibonacci
cmdh  Computer Modern Dunhill
cmssq  Computer Modern Sans Serif Quotation Style 8pt

Table 7.3: Font Families Supported by

AB AB

Figure 7.2: inch Font Examples

Besides these transparent commands for scalable fonts also offers two inch-high variants of bold Computer Modern Sans typefaces for Latin and Cyrillic: cminch and kminch. These commands bypass the NFSS font setup and should only be used for book titles etc. The command sequence {\cminch AB} {\kminch AB} produces the output shown in figure 7.2.

7.7 Shorthands for Embedding Words in a Different Typeface

Sometimes it may be necessary to give short portions of text not only in a different encoding (for which the lat{...} and mnr{...} commands are useful) but it may also be necessary to switch the typeface temporarily. Usually capsules using \textxx do the work if only the typeface is concerned, and building nested commands like \textsf{\lat{...}} is cumbersome if these changes have to be applied very often. provides an abbreviated style following the rule

[k|l]two letter font style code{...}

where the font style code is one of rm, bf, it, sl, sf, sc and tt, like

\ksl{...}, \lsc{...}, etc.

Samarin, Addison-Wesley 1994, p. 181.

Family and Command Example  Typeface Examples

Computer Modern Serif

cmr (default)

textbf{...} textsl{...} textsc{...} textit{...} fontseries{bx}\textit{...} bx

bx

>1 Computer Modern Typewriter cmtt texttt{...} texttt{\textit{...}} texttt{\textsl{...}} texttt{\textsc{...}}  Computer

Modern Typewriter example suspended…

>2 Computer Modern Variable Width Typewriter cmvtt fontfamily{cmvtt}...  cmvtt  cmvtt fontfamily{cmvtt}\textit{...}  cmvtt  cmvtt Computer

Modern Variable Width Typewriter example suspended…

>3 Computer Modern Sans cmss textsf{...} textsf{\textsl{...}} textsf{\fontseries{bx}...}  bx  bx textsf{\fontseries{sbc}...}  sbc  sbc  Computer

Modern Sans example suspended…

>4 Computer Modern Funny cmfr fontfamily{cmfr}...  cmfr  cmfr  Computer

Modern Funny example suspended… For printing the Modern Funny example the user is required to edit this file (montex.tex) at line 26 and

set the value of fontSamples to “5”.

>5 Computer Modern Dunhill cmdh fontfamily{cmdh}...  cmdh  cmdh  Computer

Modern Dunhill example suspended… For printing the Modern Dunhill example the user is required to edit this file (montex.tex) at line 26 and

set the value of fontSamples to “6”.

Table 7.4: Typeface Consistency for Cyrillic and Latin

Letter Input Letter Input

                          \ch           \Ch
                          \jh           \Jh
                          \sh           \Sh
                          \zh           \Zh
                   ŋ  \ng           \Ng
                          \g            \G

Table 7.5: Shortcuts for Mongolian Transliteration Symbols

7.8 Shorthands for Writing Transliterated Texts

provides shortcuts for writing certain accented symbols used in conventional transliterating of Mongolian by haceks, the nasal and the gamma. These shortcuts are essentially mnemonics replacing the somewhat more tedious accent notation (see table 7.5).

It must be observed that these commands are by default dependent on the environment they are used in. \Sh yields a when used in a Latin environment but results in a when used in a Cyrillic context^{5}:

agdar and adraa are transliterations for agdar and adraa.

\emph{\Sh agdar} and \emph{\Ch adraa}
are transliterations for
{\mnr\Sh agdar} and {\mnr\Ch adraa}.

7.9 Gamma Typeface

If modern Greek is supported by your L^{A}TEX 2ε installation then the shape of the gamma will match the neighbouring typeface as closely as possible=1 as can be seen from table 7.4; g otherwise, the selection of gamma shapes and styles is limited to the gamma math typeface supplied by standard TEX installations.

7.10 Oirat Double Accents

All accented characters which are contained in the T1 encoding or can be generated out of these via accents can be produced. This comes in conveniently for transliterating Oirat texts which need vowels with double diacritics, like ǟ which can be entered as any combination of two nested accent commands (like \={\"a}) or one accent command and a vowel with diacritics (provided an 8-bit input codepage is available).

5The authors wish to thank J. Knappen for resolving one instability in the original code for these letters.

7.11 Numbering by Cyrillic Letters

Analogous to the \Alpha command which provides an alphabetical counter in English, features counters for Buryat, Modern Mongolian, and Russian.

Buryat The counter for Buryat is invoked with Uzeg{n} or uzeg{n} and is

valid for 1 ≤ n ≤ 32.

11/1  ^{2}2/2  ^{3}3/3  ^{4}4/4  ^{5}5/5  ^{6}6/6
77/7  ^{8}8/8  ^{9}9/9 ^{10}10/10 ^{11}11/11 ^{12}12/12
1313/13 ^{14}14/14 ^{15}15/15 ^{16}16/16 ^{17}17/17 ^{18}18/18
1919/19 ^{20}20/20 ^{21}21/21 ^{22}22/22 ^{23}23/23 ^{24}24/24
2525/25 ^{26}26/26 ^{27}27/27 ^{28}28/28 ^{29}29/29 ^{30}30/30
3131/31 ^{32}32/32

Modern, or Xalx Mongolian The counter for Modern Mongolian is in-

voked with Useg{n} or useg{n} and is valid for 1 ≤ n ≤ 31.

11/1  ^{2}2/2  ^{3}3/3  ^{4}4/4  ^{5}5/5  ^{6}6/6
77/7  ^{8}8/8  ^{9}9/9 ^{10}10/10 ^{11}11/11 ^{12}12/12
1313/13 ^{14}14/14 ^{15}15/15 ^{16}16/16 ^{17}17/17 ^{18}18/18
1919/19 ^{20}20/20 ^{21}21/21 ^{22}22/22 ^{23}23/23 ^{24}24/24
2525/25 ^{26}26/26 ^{27}27/27 ^{28}28/28 ^{29}29/29 ^{30}30/30

3131/31

Russian The counter for Russian is invoked with Asbuk{n} or asbuk{n}

and is valid for 1 ≤ n ≤ 28.

11/1  ^{2}2/2  ^{3}3/3  ^{4}4/4  ^{5}5/5  ^{6}6/6
77/7  ^{8}8/8  ^{9}9/9 ^{10}10/10 ^{11}11/11 ^{12}12/12
1313/13 ^{14}14/14 ^{15}15/15 ^{16}16/16 ^{17}17/17 ^{18}18/18
1919/19 ^{20}20/20 ^{21}21/21 ^{22}22/22 ^{23}23/23 ^{24}24/24
2525/25 ^{26}26/26 ^{27}27/27 ^{28}28/28

Part III

Mongol Bicig and Manju Bithe munggulbicik, manjubithe.

Introduction

This part describes in detail all aspects of typesetting Mongolian and Manju with . The following sections cover the various input methods for these languages, the commands for presenting small snippets, big portions and whole documents composed in Mongolian and Manju, as well as the relationship between input notations and script-related commands.

8.1 Mongolian and Manju Script Fundamentals

Mongolian Script, or bicig, is a writing with an intriguing and complex relationship between the canonical letters of the alphabet and their presentations in context. Virtually any canonical letter can assume several shapes. As a rule of thumb, there are three or four basic shapes: the letter in isolated form, the letter in initial, medial and final position of a word. Only a few letters stay the same, and in rare cases there are up to ten possibilities for representing a single letter.

On the other hand, some letters share the same shape in different contexts; one so-called glyph can represent more than one letter, sometimes three or four different letters.

The Manju writing, or bithe system is a close relative of the Mongolian system; the basical letter shapes are the same. Yet for Manju, a set of diacritics (dots und circles) was designed to the effect that all the ambiguities of Mongolian are eliminated.

Decomposing the writing system and using glyphs as the atoms of writing is one of several conceivable methods of writing Mongolian script.

In , Mongolian script can be entered in three ways, either by writing transliterated Mongolian in one of two different romanization systems, by an approximated symbol for every glyph or by generic name. There are certain constraints concerning the possible combinations of Mongolian input methods and Mongolian writing display commands. Since Manju has only one input method, these constraints do not apply to Manju. The possible combinations are listed in table 9.2.

A complete guide to the principles of glyph analysis can be found in the MLS Report by one of the authors.^{1}

Due to technical constraints of , there is an intimate relationship between various script-related commands and Mongolian input methods.

8.2 General Settings

As for Modern (Xalx) Mongolian, Buryat and Russian documents, it is possible to set the document language to Uighur Mongolian or Manju with a language option:

\usepackage[<language options>,<encoding options>]{mls}

The two language options are bicig for Uighur Mongolian and bithe for Manju documents. Among other things, they set the document encoding, the captions and the date in either Uighur Mongolian or Manju.

The date form follows TEX conventions and is thus a mixture of numbers and words. Thus for today (July 4, 2002) we get^{2} what is shown in table 8.1.

Mongolian Manju

Figure 8.1: Dates in Uighur Mongolian and Manju

The document language option bicig can only be used with the Mongolian input method named “Simplified Transliteration” (see the following chapter and table 9.2).

1Oliver Corff: MLS Report. UNU/IIST Report No. 8, Macau 1993

2The actual date at compilation time is used for the examples.

Mongolian Input

It has been mentioned before that each Mongolian input method provided by has a slightly different application scope. The next sections cover

• Simplified Transliteration Mode which is the mode of choice for bulk text due to its availability as document encoding; the associated font encoding is labelled and internally activated by the command SetDocumentEncodingBicig.

• MLS Transliteration Mode which is most suitable for short portions of text, like dictionary entries, quotations, etc.

• Immediate Mode is the mode accepting Mongolian characters encoded in the MLS codepage. Together with this mode, the input encoding mls should be specified, ideally combined with the command SetDocumentEncodingNeutral.

• Glyph Input which is useful mainly for rendering individual words in unorthodox or incorrect spellings, e. g. for reproducing idiosyncrasies found in old books.

A comprehensive table of the Mongolian alphabet and its MLS transliteration, the input conventions of the MLS transliteration in and the Simplified Transliteration is given in table 9.1.

The possible combinations of Mongolian writing input methods and display commands are listed in table 9.2. The columns stand for each possible input encoding, the rows contain the display command types. Each table cell at the contains the command that is available for a given combination of input method and command.

9.1 Simplified Transliteration Mode

The broad romanization of the Mongolian script as realized in the MLS system focuses on lexical properties (the information layer) rather than graph-

Uighur  MLS  Simplified Uighur  MLS  Simplified
Script Transl. Input  Input  Script Transl. Input  Input

   a  a      a       a  s  s      s       s
   E  ä     ä, E      e  S  sh      S       sh
   e  e      e       v  t  t      t       t
    i  i      i       i  d  d      d      d, t
   o  o      o       u  l  l      l       l
   u  u      u       u  m  m      m       m
   O  ö     ö, O    ui, u  c  c      c       c
   U  ü     ü, U    ui, u  z  z      z       z
   n  n      n       n  y  y      y       y
   |ng  *ng     ng      ng  r  r      r       r
   x  x      x       x  v  v      v       v
   G  G       g  h  h      h       h
   k  k      k       k  j  j      j       j
   g  g      g      g, k  K  K      K       K
   b  b      b       b  Q  [–]     Q       q
   p  p      p       p  C  C      C       C
    f  f      f       f  Z  Z      Z       Z

Table 9.1: Mongolian Script Transliterations

ical properties (the presentation layer). The obvious advantage of such a method is the possibility to store and transmit Mongolian language information in and between systems without devices for displaying Mongolian writing.

With the ambiguities of the Mongolian script (the consonants t/d, the vowels a/e, and many other ambiguous shapes give vivid evidence hereof) it is however possible to enter misleading or wrong romanizations which lead to a desired yet semantically misleading display of Mongolian in which case the underlying information is not suitable for further processing.

Another aspect is the retrieval of information from e. g. library catalogues when only the display of potentially unknown words like in book titles is available. It must be possible to enter Mongolian script into an information processing system without knowing at every moment which underlying letter generates a given shape. This implies that glyph analysis does not decompose complex glyph shapes into atoms if the shape transformation is purely dictated by graphical rather than linguistical context.

Notwithstanding this fact, an obvious n appearing as following a vowel should be entered as n while an a following a consonant which also appears as should certainly entered as vowel, not as consonant. Anything going deeper in glyph analysis can only be considered as atomic coding which may be highly useful in special cases but renders the input process more than

Command  Mongolian  Manju
Type  MLS  Simplified

Document Encoding only available as font encoding , not as document encoding LMO

LMA

Horizontal Capsules bcg bicig

bithe

Horizontal Paragraphs not available bicigtext

bithetext

Vertical Capsules mbosoo mobosoo

mabosoo

Vertical Paragraph Boxes not available mobox

mabox

Vertical Pages not available bicigpage

bithepage

Table 9.2: Mongolian Input and Display Commands

cumbersome in general cases.

The Mongolian Simplified Transliteration proposed here is based on principles laid out by Dr. Michael Balk of the Deutsche Staatsbibliothek, Stiftung Preussischer Kulturbesitz, Berlin. During its development, various proposals were discussed at DIN, MNISM and ISO standardization meetings during 1994 to 1997.

The most important principle of this simplified input method is the consequent elimination of ambiguities in the relation between romanized input (as performed on an ordinary computer keyboard) and its Mongolian script target. If the Mongolian language provides several readings for certain vowels, then only one vowel is available in the simplified method; if alternating consonants (like k/g can swap shapes, then each input letter is associated with one and only one output shape. Furthermore, the input alphabet (speaking in terms of computer theory) is limited to the basic Latin alphabet. It uses only

a b c d e f g h i j k l m n p q r s t u v x y z C K Z and the fol-

lowing characters with special meanings: - = ' ". The first symbol is used to separate grammatical endings from preceding words, the second separates floating vowels from word stems, and the third and fourth character act as Variant Selectors. At present, the second Variant Selector is not yet assigned.

Unlike a purely atomic rendering, the resulting romanization as shown in table 9.1 is easy to learn, much easier to read than atomic code, and yet acceptably close to conventional Mongolian transliterations, as can be seen from table 9.3. Each row contains one or more instances of every letter listed in the first column. From left to right, these are the beginning of a word, the middle of a word and the end of a word. Every single cell features three elements: the example in MLS romanization appears in italics; the correct Simplified Transliteration input appears in typewriter style, and the word in Uighur Mongolian letters appears in the right half of the cell.

Letter  Beginning  Middle  End
a  arad arad  arad  baatur bagadur  bagadur {la \atop la}  la
sana=a sanag=a  sanag=a
ä  ärkä erke  erke  cäcäg cecek  cecek  sükä suike  suike
e  {eKs \atop evKs}  evKs  geologi kvuluki  kvuluki
i  izaur izagur  izagur  minu minu  minu  bandi bandi  bandi
o  olan ulan  ulan  aul=a agul=a  agul=a
u  ulus ulus  ulus  arxu garxu  garxu
ö  öndür uindur  uindur  cöm=ä cuim=e  cuim=e
ü  üsüg uisuk  uisuk
n + \{V\} {nam \atop nam}  nam  onol unul  unul  bayin=a baiin=a  baiin=a
n + {C}  bandi bandi  bandi
n’ + \{V\} {n’am \atop n'am}  n’am  on’ol un'ul  un’ul  bayin’=a baiin'=a  baiin’=a
n’ + {C}  KoNTor Kun'tur  Kun’tur ban’di ban'di  ban’di
ng  mongol munggul  munggul vang vang  vang
x  xota xuda  xuda  abxu abxu  abxu  mix=a mix=a  mix=a
azar gazar  gazar  baatur bagadur  bagadur {tu \atop tug}  tug
’  ’azar g'azar  g’azar  ba’atur bag'adur  bag’adur

k  käräg kerek  kerek  ärkiläkü erkileku  erkileku
g  {gär \atop ger}  ger  ügäi uigei  uigei  bicig bicik  bicik
b  {ba=a \atop bag=a}  bag=a  däbtär tebder  tebder  {äb \atop eb}  eb
p  pangsa pangsa  pangsa
f  feodal fvudal  fvudal  Cifr Cifr  Cifr
s  saxal saxal  saxal  basa basa  basa  {nas \atop nas}  nas
sh  shadur shagdur  shagdur
t  tomu tumu  tumu  baatur bagadur  bagadur
d  dumdadu dumdadu  dumdadu {odu \atop udu}  udu  arad arad  arad
sädgil sedkil  sedkil  {äD \atop ed'}  ed’
l  {la \atop la}  la  aldar aldar  aldar  onul unul  unul
blam=a blam=a  blam=a
m  mongol munggul  munggul nomin numin  numin  {nom \atop num}  num
c  caan cagan  cagan  äcän ecen  ecen
z  {zam \atop zam}  zam  azar gazar  gazar
y  yondan yundan  yundan  bayar bayar  bayar  xoriy=a xuriy=a  xuriy=a
r  rashan rashan  rashan  oros urus  urus  bolor bulur  bulur
v  vang vang  vang
h  heze hvzv  hvzv  lhas lhas  lhas
j  {j \atop j}  j
K  KoNTor Kun'tur  Kun’tur

gh  ghombo qumbu  qumbu
C  Cifr Cifr  Cifr  sTan’C stan'C  stan’C
Z  Zambu Zambu  Zambu  {aZi \atop aZi}  aZi

Table 9.3: Mongolian Simplified Transliteration by Example

While the input method for the majority of characters matches the transliteration conventions, some letters require a slightly different treatment:

1. Although the diphtong *aii* is usually rendered as ayi, it must be entered as aii in order to produce the desired effect.

2. The back vowels o and u are both rendered as u.

3. The front vowels ö and ü are both rendered as ui in first syllables and as u in later syllables.

4. Since t means both t and d, it is necessary to spell this letter as t in the beginning of words, and d in the middle of words, regardless of the actual meaning.

5. The four consonants , g, x and k are constrained with regard to the following vowels. The Simplified Transliteration renders these as g (before a and u only), g (before a and u only), x and k.

As it was demonstrated in section 7.1, it is technically possible to choose between an automatic document encoding and the neutral mode. In the case of Uighur Mongolian, the mode of choice activates the Simplified Transliteration Mode and is called with

     \SetDocumentEncodingBicig

With \SetDocumentEncodingBicig set, it is possible to switch to the Simplified Transliteration Mode anywhere in the document, not only in the preamble.

Caveat: Since switching to Uighur Mongolian text requires a lot of settings to be effected at the same time, there are high-level commands available (see below, chapter ??: Mongolian and Manju Display Commands) which do all the work, including the definition of the document encoding. Thus, while \SetDocumentEncodingBicig is indeed classified as a user-level command, it is certainly not necessary for everyday work.

9.1.1 Character Variants

With the assistance of special, non-printing characters like the Form Variant Selectors, the appearance of certain characters can be modified in order to display typographical and orthographical variants. Notably, the n will loose its dot before vowels, as will . Let’s assume the word “place” is written in an old book as g’azar. It should be understood that this is a variant of gazar and should be spelled ’azar, not xazar. With vowels, the Form Variant Selectors can change the shape that is usually required by graphical context. At present, only the first of two Form Variant Selectors actually does something, the exact behaviour of the second Form Variant Selector waits to be implemented.

The following short example shows a concrete application of this method.

It renders the six syllable mantra om ma ni padme hum (tib. �� ��� � ��

��{� � \atop ��}��{� \atop {� \atop � �} ��} ��) also featuring the special syllable om as it is displayed on a huge

bronze incense burner in front of the Gandan Monastery in Ulaanbaatar:

3cm�uva ma’=a n’i badmi’

�huu

\mobox{3cm}{\noindent\sffamily
\om uva\\
\  ma'=a\\
\  n'i\\
\  badmi'\\
\om huu}

Figure 9.1: Mongolian Character Variants Example

9.2 MLS Transliteration Mode

In Transliteration Mode (activated with the commands bcg{...} or mbosoo{...}) Mongolian text portions can be entered using a transliteration which is a rough approximation to the MLS system. The major difference is that only pure Latin alphabetical symbols can be used for virtually all letters. Front vowels are either entered via the traditional vowels with diacritics (ä, ö, ü) or can be entered with capitalized versions of the normal